# Beyond RMSE and MAE: Introducing EAUC to unmask hidden bias and unfairness in dyadic regression models

### v3

@@ Abstract
Dyadic regression models, which output real-valued predictions for pairs of entities, are fundamental in many domains (e.g. obtaining user-product ratings in Recommender Systems) and promising and under exploration in others (e.g. tuning patient-drug dosages in personalized pharmacology). In this work, we prove that non-uniform observed value distributions of individual entities lead to severe biases in state-of-the-art models, skewing predictions towards the average of observed past values for the entity and providing worse-than-random predictive power in eccentric yet crucial cases; we name this phenomenon eccentricity bias. We show that global error metrics like Root Mean Squared Error (RMSE) are insufficient to capture this bias, and we introduce Eccentricity-Area Under the Curve (EAUC) as a novel complementary metric that can quantify it in all studied domains and models. We prove the intuitive interpretation of EAUC by experimenting with naive post-training bias corrections, and theorize other options to use EAUC to guide the construction of fair models. This work contributes a bias-aware evaluation of dyadic regression to prevent unfairness in critical real-world applications of such systems.
@@ II-A Defining Dyadic Data and Dyadic Regression
Dyadic data consists of interactions between pairs of entities, where each observation involves a user interacting with an item. Tasks on this data typically involve performing a prediction given a pair of entities (e.g. predicting whether a job applicant and a vacant position are compatible) or predicting the most optimal pair of entities given the task requirements (e.g. predicting the most suitable applicant for a job position).
@@
This work focuses on dyadic regression tasks, which require predicting a real value for any dyad by determining the optimal regressor function $f$:
$f:\mathcal{U}\times\mathcal{I}\rightarrow\mathbb{R}$
such that $f$ predicts the optimal $\hat{r}_{ui}\in\mathbb{R}$ for dyad $(u,i)$.
@@ II-B State-of-the-art evaluation: RMSE and Netflix Prize
While RMSE is simple to interpret (lower values indicate lower overall prediction error), facilitating its use and that of similar metrics in the field. Other popular datasets like MovieLens [15] further cemented global error metrics as the sole standard for evaluating dyadic regression tasks [16, 17]. However, they only offer an image of the overall error of the regressor, meaning they can potentially neglect biases in the predictions.
@@ III Motivation: Limitations of RMSE in Dyadic regression tasks
Table I shows the Test RMSE for the examples of each unique rating (1 to 5) and the entire Test set. Although the system performs best for reviews with ratings 3 or 4, the most common in the data ($\sim 28\%$ and $\sim 33\%$ of ratings, respectively), predictions are always notably better than random, suggesting the absence of relevant biases causing unfairness.
@@
where $I_{u}$ is the set of items that $u$ interacted with, $U_{i}$ is the set of users who interacted with item $i$, and $r_{ui}$ is an observed value.
@@
To observe this condition from an alternative perspective, Figure 3 shows the relation between ratings and predictions for two different subsets of users and items: test cases $(u,i)$ where $2.0\leq DMV_{ui}\leq 2.5$ (user-item combinations with low average observed values), and test cases with $4.0\leq DMV_{ui}\leq 4.5$ (user-item combinations with high average observed values). As seen in Table I, MF is mostly unbiased if viewing the dataset as a whole, but in the low and high $DMV_{ui}$ specific scenarios we observe that the predictions are severely skewed towards the $DMV_{ui}$ of the dyad: the model assumes a critical user will not ever like a low-rated film even if they gave it a high rating, and vice-versa: we name this phenomenon as eccentricity bias. As such, the knowledge generalization of the model is lower than an RMSE-guided evaluation would suggest: most of the global error minimization is achieved by blindly skewing predictions towards average observed past values, rather than by learning latent user and item characteristics.
@@
IV Bias-aware Evaluation of Dyadic Regression Models
@@ IV-A Eccentricity-Area Under the Curve (EAUC)
    Sort Test examples in ascending order of eccentricity, creating an ordered sequence of pairs $(Ecc_{j},\epsilon_{j})$ with increasing $Ecc_{j}$ values.
@@
Inputs:
  • Train examples $D_{\text{train}}=\{(u,i,r_{ui})\}$
  • Test examples $D_{\text{test}}=\{(u,i,r_{ui}$
  • ($u\in\mathcal{U},i\in\mathcal{I},r_{ui}\in\mathbb{R},\hat{r}_{ui}\in\mathbb{R}$)
@@ IV-B Interpretation of EAUC for model evaluation
Consequently, EAUC offers a valuable complement to traditional metrics like RMSE and MAE, which quantify average errors but neglect predictive bias learnt by models to optimise their predictions in dyadic settings. Moreover, since eccentricity bias leads to discriminatory outcomes on tasks with strict fairness requirements, EAUC also behaves as measure of fairness in dyadic regression tasks.
@@ V-A Datasets
We utilize dyadic regression datasets, where each observation is a real-valued interaction between two entities, in three contexts with significant fairness implications: biomedicine and pharmacology, product recommendation, and finances and lending. Table II shows basic statistics of the datasets selected for our experiments: Netflix Prize [11] and Movielens 1M [15] are product recommendation datasets, where the objective is to predict the rating a given user will give to a movie after viewing it. GDSC1 [22] and CTRPv2 [23] contain biomedical pharmacology data, where the goal is to predict the effectiveness of a given drug in fighting against specific cancer cell lines. IMF DOTS 2023 [24] and Kiva [25] contain financial interaction data; the former seeks to predict the potential Value of Trade between two economic entities (in this case, UN countries or regions), and the later consists in predicting the normalized lent amount by a lender to a specific cluster of microloans.
TABLE II: Basic statistics of the used datasets.
Domain Dataset Examples Users Items Density Product Recommendation Movielens 1M 1,000,209 6,040 3,706 4.5% Netflix Prize 100,480,507 480,189 17,770 1.2% Biomedical & Pharmacology GDSC1 99,948 969 402 25.6% CTRPv2 108,224 829 544 24.0% Financial & Lending IMF DOTS 2023 16,776 233 233 32.0% Kiva Microloans 110,371 4,005 2,673 1.0%
@@
We applied 90/10 random train/test data splitting to all datasets, and used only implicit feedback information (no extra features). For GDSC1, CTRPv2 and IMF DOTS 2023, we sampled $1/3$ of the total examples before performing the train-test split, to reproduce a higher sparsity more similar to that of real-world use-case scenarios of dyadic regression.
@@ VI Results and Discussion
Movielens 1M $D_{KS_{\mathcal{D}}}$: 0.412 Netflix Prize $D_{KS_{\mathcal{D}}}$: 0.429 RMSE MAE EAUC RMSE MAE EAUC Random 1.710 ± 0.004 1.397 ± 0.004 0.361 ± 0.016 1.696 ± 0.000 1.386 ± 0.001 0.395 ± 0.004 Dyad Average 0.956 ± 0.002 0.770 ± 0.002 0.370 ± 0.013 0.964 ± 0.000 0.781 ± 0.000 0.435 ± 0.008 Bayesian SVD++ 0.830 ± 0.003 0.650 ± 0.002 0.297 ± 0.013 - - - Glocal-K 0.826 ± 0.003 0.643 ± 0.002 0.297 ± 0.029 - - - GC-MC 0.846 ± 0.005 0.665 ± 0.004 0.323 ± 0.028 - - - MF 0.840 ± 0.003 0.657 ± 0.003 0.315 ± 0.013 0.826 ± 0.001 0.635 ± 0.001 0.354 ± 0.006
@@
GDSC1 $D_{KS_{\mathcal{D}}}$: 0.384 CTRPv2 $D_{KS_{\mathcal{D}}}$: 0.390 RMSE MAE EAUC RMSE MAE EAUC Random 6.964 ± 0.069 5.864 ± 0.066 0.134 ± 0.015 5.804 ± 2.297 5.689 ± 0.072 0.207 ± 0.036 Dyad Average 1.672 ± 0.016 1.301 ± 0.010 0.081 ± 0.008 1.841 ± 0.011 1.332 ± 0.007 0.179 ± 0.016 Bayesian SVD++ 1.029 ± 0.035 0.747 ± 0.007 0.039 ± 0.006 1.244 ± 0.010 0.865 ± 0.006 0.074 ± 0.011 Glocal-K 1.063 ± 0.020 0.780 ± 0.015 0.048 ± 0.007 1.241 ± 0.014 0.866 ± 0.007 0.086 ± 0.014 GC-MC 1.041 ± 0.008 0.754 ± 0.007 0.047 ± 0.005 1.220 ± 0.019 0.841 ± 0.011 0.078 ± 0.020 MF 1.074 ± 0.015 0.795 ± 0.011 0.042 ± 0.005 1.279 ± 0.020 0.902 ± 0.006 0.105 ± 0.012
@@
IMF DOTS 2023 $D_{KS_{\mathcal{D}}}$: 0.3717 Kiva Microloans $D_{KS_{\mathcal{D}}}$: 0.873 RMSE MAE EAUC RMSE MAE EAUC Random 4.044 ± 0.102 3.350 ± 0.104 0.189 ± 0.018 $1$ 0.003 0.952 ± 0.004 0.408 ± 0.006 Dyad Average 1.355 ± 0.033 1.076 ± 0.034 0.136 ± 0.019 0.413 ± 0.009 0.227 ± 0.004 0.498 ± 0.003 Bayesian SVD++ 0.937 ± 0.018 0.664 ± 0.013 0.080 ± 0.024 0.472 ± 0.007 0.258 ± 0.003 0.461 ± 0.011 Glocal-K 0.889 ± 0.013 0.640 ± 0.014 0.075 ± 0.015 0.415 ± 0.012 0.213 ± 0.003 0.461 ± 0.020 GC-MC 0.946 ± 0.049 0.658 ± 0.029 0.080 ± 0.021 0.414 ± 0.009 0.226 ± 0.009 0.469 ± 0.015 MF 1.005 ± 0.012 2.111 ± 0.046 0.100 ± 0.017 0.405 ± 0.009 0.205 ± 0.005 0.440 ± 0.007
@@
Figure 5 shows for each dataset the relation between the eccentricity of observed values and the prediction errors. Model errors are near-zero when the eccentricity is very small but increase almost linearly with eccentricity. This is particularly severe in datasets with very non-uniform observed value distributions, such as Netflix and Kiva, in which a global optimizer will be rewarded for predicting always values near $DMV_{ui}$. Indeed, it can also be observed in Table III that the measured $D_{KS_{\mathcal{D}}}$ for a dataset correlates with the measured EAUC, indicating that less uniform local distributions of observed values of users and items lead to learnt models with more severe eccentricity biases.
Product Recommendation
Biomedicine & Pharmacology
Financial & Lending
@@
Summarizing, these results prove that the evaluation of dyadic regression tasks should not be limited to global error metrics like RMSE or MAE; instead, it should also consider the evaluation of entity-wise biases, which we achieve with our novel metric EAUC.
@@ VI-A Usage of EAUC to mitigate eccentricity bias
As we proved EAUC can be used as a reliable metric to measure the discovered eccentricity bias and contribute a fairness-aware evaluation of dyadic regression models, we theorize that EAUC itself can be used to aid in obtaining fair dyadic regression models which mitigate or avoid eccentricity bias. While the development and evaluation of such models is out of the scope of this article, Figure 7 suggests some potential avenues for fairness-oriented training. One naive option (Subfigure a) is to derive correction functions based on eccentricity for de-biasing the unfair predictions of regression models trained with global error optimizers; the usage of such fairness-oriented corrections has seen success in different contexts, such as gender-fair top-N recommender systems [29]. Another potential approach (Subfigure b) would be to include EAUC as a fairness constraint in the model’s optimizer during training as a penalty term [30] or through adversarial learning [31]. Finally, EAUC could be integrated as the fairness component of evaluation criteria in multi-objective Auto-ML frameworks (Subfigure c) to enable fairness-aware model selection [32].
@@
Table IV and Figure 8 show the effects of the corrections in the global error (RMSE, MAE) and bias-aware performance (EAUC) and its curve representation in an MF model and the Test set of Netflix Prize. Although they should not be interpreted as a foolproof de-biasing prediction, their effects are evident. While the corrections obtain slightly higher RMSE than the original model on examples with near-zero eccentricities, they maintain a much more restrained error as eccentricity increases; this, in turn, translates to a lower EAUC value obtained by the correction, corresponding to lower eccentricity bias and less unfairness in contexts with such implications. Overall, this corroborates the adequacy of EAUC and its potential applicability to guide the construction of fair models.
@@
RMSE MAE EAUC Random 1.696 ± 0.000 1.386 ± 0.001 0.395 ± 0.004 Dyad Average 0.964 ± 0.000 0.781 ± 0.000 0.435 ± 0.008 MF 0.826 ± 0.001 0.635 ± 0.001 0.354 ± 0.006 w/ Linear corr. 1.186 ± 0.004 0.894 ± 0.003 0.253 ± 0.001 w/ Linear corr. + ML-RUS (Clipping) 0.879 ± 0.001 0.686 ± 0.001 0.307 ± 0.001 w/ Linear corr. + ML-RUS (Sigmoid) 0.977 ± 0.001 0.784 ± 0.002 0.267 ± 0.001 w/ Random Forest corr. 1.008 ± 0.001 0.809 ± 0.001 0.276 ± 0.001

### v4

@@ Abstract
Dyadic regression models, which output real-valued predictions for pairs of entities, are fundamental in many domains (e.g. obtaining user-product ratings in Recommender Systems) and promising and under exploration in others (e.g. tuning patient-drug dosages in precision pharmacology). In this work, we prove that non-uniform observed value distributions of individual entities lead to severe biases in state-of-the-art models, skewing predictions towards the average of observed past values for the entity and providing worse-than-random predictive power in eccentric yet crucial cases; we name this phenomenon eccentricity bias. We show that global error metrics like Root Mean Squared Error (RMSE) are insufficient to capture this bias, and we introduce Eccentricity-Area Under the Curve (EAUC) as a novel metric that can quantify it in all studied domains and models. We prove the intuitive interpretation of EAUC by experimenting with naive post-training bias corrections, and theorize other options to use EAUC to guide the construction of fair models. This work contributes a bias-aware evaluation of dyadic regression to prevent unfairness in critical real-world applications of such systems.
@@ II-A Defining Dyadic Data and Dyadic Regression
Dyadic data comprises interactions between pairs of entities, where each observation depicts a user interacting with an item. Typical tasks involve predicting a value given a pair of entities (e.g. predicting whether a job applicant and a vacant position are compatible) or predicting the optimal pair of entities given the task requirements (e.g. predicting the most suitable applicant for a job position).
@@
This work focuses on dyadic regression tasks, which require predicting a real value for any dyad by determining the optimal regressor function $f:\mathcal{U}\times\mathcal{I}\rightarrow\mathbb{R}$ such that $f$ predicts the optimal $\hat{r}_{ui}\in\mathbb{R}$ for dyad $(u,i)$.
@@ II-B State-of-the-art evaluation: RMSE and Netflix Prize
RMSE is simple to interpret (lower is better), facilitating its use. Other popular datasets like MovieLens [15] further cemented RMSE and other global error metrics as the sole standard for evaluating dyadic regression tasks [16, 17]. However, they only offer an image of the overall error of the regressor,potentially neglecting biases in the predictions.
@@ III Motivation: Limitations of RMSE in Dyadic regression tasks
Table I shows the Test RMSE for the examples of each unique rating (1 to 5) and the entire Test set. Although the system performs best for reviews with ratings 3 or 4 (the most common with $\sim 28\%$ and $\sim 33\%$ of ratings, respectively), predictions are always notably better than random, suggesting the absence of unfairness due to bias.
@@
where $I_{u}$ and $U_{i}$ are the set of items and users that $u$ and $i$ interacted with, respectively, and $r_{ui}$ is an observed value.
@@
To observe this condition from an alternative perspective, Figure 3 shows the relation between ratings and predictions in two scenarios: test cases $(u,i)$ where $2.0\leq DMV_{ui}\leq 2.5$ (user-item combinations with low average observed values), and test cases with $4.0\leq DMV_{ui}\leq 4.5$ (user-item combinations with high average observed values).
@@
Although MF is mostly unbiased if viewing the dataset as a whole, in the low and high $DMV_{ui}$ scenarios we observe that the predictions are severely skewed towards the $DMV_{ui}$ of the dyad: the model assumes a critical user will not like a low-rated film even if they gave it a high rating, and vice-versa: we name this phenomenon as eccentricity bias. As such, the knowledge generalization of the model is lower than an RMSE-guided evaluation would suggest: most of the global error minimization is achieved by blindly skewing predictions towards average observed values rather than by learning latent user and item features.
@@
IV Bias-aware Evaluation of Dyadic Regression
@@ IV-A Eccentricity-Area Under the Curve (EAUC)
    Sort Test examples into an ordered sequence of pairs $(Ecc_{j},\epsilon_{j})$ with increasing $Ecc_{j}$ values.
@@
Inputs: Train and test examples $D_{\text{train}}=\{(u,i,r_{ui})\}$, $D_{\text{test}}=\{(u,i,r_{ui})\}$
@@
The EAUC of a given model $f:\mathcal{U}\times\mathcal{I}\rightarrow\mathbb{R}$ in a given dataset $\mathcal{D}$ with training and test subsets $\mathcal{D}_{train}$ and $\mathcal{D}_{test}$ may also be articulated as a differentiable function $EAUC(f,\mathcal{D})$ in formulaic fashion, such that:
$EAUC(f,\mathcal{D})=\frac{\sum\limits_{x_{i}\in\mathcal{D}_{test}}\frac{(Ecc_{x_{i}}-Ecc_{x_{i-1}})*(\epsilon_{x_{i}}+\epsilon_{x_{i-1}})}{2}}{(max(r_{ui})-min(r_{ui}))^{2}}\quad$ (5) $\text{Where}\quad\quad\quad Ecc_{\{u,i,r_{ui}\}}=$ $|r_{ui}-\frac{1}{2}(\bar{r}_{u}+\bar{r}_{i})|,$ $\epsilon_{\{u,i,r_{ui}\}}=$ $(r_{ui}-f(u,i))^{2},$ $\bar{r}_{u}=$ $\frac{1}{|\mathcal{I}_{u}|}\sum_{i\in\mathcal{I}_{u}\subseteq\mathcal{I}_{\mathcal{D}_{train}}}r_{ui},$ $\bar{r}_{i}=$ $\frac{1}{|\mathcal{U}_{i}|}\sum_{u\in\mathcal{U}_{i}\subseteq\mathcal{U}_{\mathcal{D}_{train}}}r_{ui},$ $x_{i}\in$$\{u,i,r_{ui}\};$ $\text{With}\quad(Ecc_{x_{i}}-Ecc_{x_{i-1}})\geq$ 00 0
@@ IV-B Interpretation of EAUC for model evaluation
Consequently, EAUC offers a valuable complement to traditional metrics like RMSE and MAE, which quantify average errors but neglect predictive bias learnt by models to optimise their predictions in dyadic settings.
Moreover, since eccentricity bias leads to discriminatory outcomes when fairness is involved, EAUC also behaves as measure of fairness in dyadic regression tasks, being able to complement and synergize with classic fairness metrics. In this aspect, it is plausible that EAUC will become a reliable detector of an -until now- unknown source of unfairness in this type of task, i.e. eccentricity bias. As an indicator of unfairness in dyadic regression, EAUC may be used to directly detect the presence of eccentricity-based individual or group unfairness [19], which can later be characterized using utility or parity-based fairness metrics [20]. Conversely, the inverse is also possible: while most group or individual unfairness are often attributed to representational or class-stereotypical biases in the data [21], they often cannot fully explain the source of fairness[22]; EAUC and its analysis may be used to better characterize the source of otherwise opaque unfairness phenomenons. One final property of EUAC is its ability to, unlike classic individual and group-based metrics modelling parity and utility [23], detect eccentricity-based unfairness that affects all entities (e.g. users) of the population as a whole, rather than only specific subgroups or individuals. As a result, EAUC can improve our ability to both detect and better characterize new and known unfairness in dyadic regression.
@@ V-A Datasets
We selected for our experiments dyadic regression datasets in three contexts with significant fairness implications: biomedicine and pharmacology, product recommendation, and finances and lending.
Table II shows basic statistics of the used datasets: Netflix Prize [11] and Movielens 1M [15] are product recommendation datasets, where the objective is to predict the rating a given user will give to a movie after viewing it. GDSC1 [27] and CTRPv2 [28] contain biomedical pharmacology data, where the goal is to predict the effectiveness of a given drug in fighting against specific cancer cell lines. IMF DOTS 2023 [29] and Kiva [30] contain financial interaction data; the former seeks to predict the potential Value of Trade between two economic entities (in this case, UN countries or regions), and the later consists in predicting the normalized lent amount by a lender to a specific cluster of microloans.
TABLE II: Basic statistics of the used datasets
Domain Dataset Examples Users Items Density
Product Recommendation Movielens 1M 1,000,209 6,040 3,706 4.5%
Netflix Prize 100,480,507 480,189 17,770 1.2%
Biomedical & Pharmacology GDSC1 99,948 969 402 25.6%
CTRPv2 108,224 829 544 24.0%
Financial & Lending IMF DOTS 2023 16,776 233 233 32.0%
Kiva Microloans 110,371 4,005 2,673 1.0%
@@
We applied 90/10 random train/test data splitting to all datasets, and used only implicit feedback information (no extra features). For GDSC1, CTRPv2 and IMF DOTS 2023, we sampled $1/3$ of the data before performing the train-test split, to reproduce a higher sparsity more akin to real-world use-case scenarios of dyadic regression.
@@ VI Results and Discussion
Movielens 1M $D_{KS_{\mathcal{D}}}$: 0.412 Netflix Prize $D_{KS_{\mathcal{D}}}$: 0.429
RMSE MAE EAUC RMSE MAE EAUC
Random 1.710 ± 0.004 1.397 ± 0.004 0.361 ± 0.016 1.696 ± 0.000 1.386 ± 0.001 0.395 ± 0.004
Dyad Average 0.956 ± 0.002 0.770 ± 0.002 0.370 ± 0.013 0.964 ± 0.000 0.781 ± 0.000 0.435 ± 0.008
Bayesian SVD++ 0.830 ± 0.003 0.650 ± 0.002 0.297 ± 0.013 - - -
Glocal-K 0.826 ± 0.003 0.643 ± 0.002 0.297 ± 0.029 - - -
GC-MC 0.846 ± 0.005 0.665 ± 0.004 0.323 ± 0.028 - - -
MF 0.840 ± 0.003 0.657 ± 0.003 0.315 ± 0.013 0.826 ± 0.001 0.635 ± 0.001 0.354 ± 0.006
@@
GDSC1 $D_{KS_{\mathcal{D}}}$: 0.384 CTRPv2 $D_{KS_{\mathcal{D}}}$: 0.390
RMSE MAE EAUC RMSE MAE EAUC
Random 6.964 ± 0.069 5.864 ± 0.066 0.134 ± 0.015 5.804 ± 2.297 5.689 ± 0.072 0.207 ± 0.036
Dyad Average 1.672 ± 0.016 1.301 ± 0.010 0.081 ± 0.008 1.841 ± 0.011 1.332 ± 0.007 0.179 ± 0.016
Bayesian SVD++ 1.029 ± 0.035 0.747 ± 0.007 0.039 ± 0.006 1.244 ± 0.010 0.865 ± 0.006 0.074 ± 0.011
Glocal-K 1.063 ± 0.020 0.780 ± 0.015 0.048 ± 0.007 1.241 ± 0.014 0.866 ± 0.007 0.086 ± 0.014
GC-MC 1.041 ± 0.008 0.754 ± 0.007 0.047 ± 0.005 1.220 ± 0.019 0.841 ± 0.011 0.078 ± 0.020
MF 1.074 ± 0.015 0.795 ± 0.011 0.042 ± 0.005 1.279 ± 0.020 0.902 ± 0.006 0.105 ± 0.012
@@
IMF DOTS 2023 $D_{KS_{\mathcal{D}}}$: 0.3717 Kiva Microloans $D_{KS_{\mathcal{D}}}$: 0.873
RMSE MAE EAUC RMSE MAE EAUC
Random 4.044 ± 0.102 3.350 ± 0.104 0.189 ± 0.018 $1$ 0.003 0.952 ± 0.004 0.408 ± 0.006
Dyad Average 1.355 ± 0.033 1.076 ± 0.034 0.136 ± 0.019 0.413 ± 0.009 0.227 ± 0.004 0.498 ± 0.003
Bayesian SVD++ 0.937 ± 0.018 0.664 ± 0.013 0.080 ± 0.024 0.472 ± 0.007 0.258 ± 0.003 0.461 ± 0.011
Glocal-K 0.889 ± 0.013 0.640 ± 0.014 0.075 ± 0.015 0.415 ± 0.012 0.213 ± 0.003 0.461 ± 0.020
GC-MC 0.946 ± 0.049 0.658 ± 0.029 0.080 ± 0.021 0.414 ± 0.009 0.226 ± 0.009 0.469 ± 0.015
MF 1.005 ± 0.012 2.111 ± 0.046 0.100 ± 0.017 0.405 ± 0.009 0.205 ± 0.005 0.440 ± 0.007
@@
Figure 5 shows for each dataset the relation between the eccentricity of observed values and the prediction errors. Model error is near-zero when the eccentricity is very low but increases almost linearly with eccentricity. This is particularly severe in datasets with very non-uniform observed value distributions, such as Netflix and Kiva, whereglobal optimizers will be rewarded for always predicting values near $DMV_{ui}$. Indeed, it can also be observed in Table III that the $D_{KS_{\mathcal{D}}}$ for a dataset correlates with the measured EAUC, indicating that less uniform local distributions of observed values of users and items lead to learnt models with more severe eccentricity biases.
Product Recommendation Refer to caption Refer to caption
Biomedicine & Pharmacology Refer to caption Refer to caption
Financial & Lending Refer to caption Refer to caption
@@
Summarizing, we have proven the evaluation of dyadic regression should not be limited to global error metrics like RMSE; it should also consider the evaluation of entity-wise biases, which we achieve with our novel metric EAUC.
@@ VI-A Usage of EAUC to mitigate eccentricity bias
Until now, we proved EAUC can be used as a reliable metric to detect and quantify eccentricity bias and contribute to a fairness-aware evaluation of dyadic regression; we theorize that EAUC can also guide the obtaining of fair dyadic regression models which mitigate or avoid eccentricity bias. While the development of such models is out of the scope of this article, Figure 7 suggests some potential avenues for fairness-oriented training and selection:
  1. (a)
    One naive option is to derive eccentricity-based correction functions to de-bias the unfair predictions of biased trained models. This is, for a biased dyadic regression model $f$ we could define a correcting $\phi$ that remaps the outputs of $f$ such that:
    $\hat{r_{ui}}_{corr}=\phi(f(u,i),\dots)$
    $EAUC(\phi\circ f,\mathcal{D})<EAUC(f,\mathcal{D})$ (7)
    Post-training bias-correction methods have the benefit of taking advantage of the already trained model, avoiding the computational complexity of training a fair model from scratch (as the correction function $\phi$ will be typically simpler). In fairness contexts, a similar approach has been previously used, for instance, to correct gender-based demographic biases in Recommender Systems [34, 35].
  2. (b)
    Another possibility would be to directly include EAUC as a fairness constraint in the dyadic regressor’s optimizer during training; an interesting property of EAUC is that, unlike most AUC-based metrics, it is differentiable for any $\mathcal{D}$ with $|\mathcal{D}|\geq 2$. Therefore, we could define a combined loss $\mathcal{L}_{fair}$ such that:
    $\mathcal{L}_{fair}(B)$ $=\mathcal{L}_{MSE}(B)+\mathcal{L}_{EAUC}(B)$
    $=\frac{\sum\limits_{(u,i,r_{ui})\in B}(r_{ui}-f(u,i))^{2}}{|B|}+EAUC(f,B)$ (8)
    where $B\subseteq\mathcal{D}_{train}$ is a training batch. Similar one-step fair training paradigms have been used with more indirect parity and utility metrics by Yao and Huang [20] to mitigate gender bias in movie recommendation, or through compositional adversarial learning by Bose and Hamilton [36], also in recommendation contexts.
  3. (c)
    Yet another alternative would be to integrate EAUC as the fairness component of evaluation in multi-objective Auto-ML frameworks; recently, Freitas [37] proposed a conceptual framework based on the combination of Pareto and lexicographic approaches, where EAUC could be used as the main fairness criterion.
@@
Table IV and Figure 8 show the effects of the corrections in the global error (RMSE, MAE) and eccentricity bias (EAUC) and its curve analysis in an MF model and the Test set of Netflix Prize. Although they should not be interpreted as foolproof de-biasing corrections, their effects are evident. While they carry slightly higher RMSE than the original model on near-zero eccentricities, they maintain a much more restrained error as eccentricity increases; this, in turn, translates to a lower EAUC obtained by the correction, corresponding to lower eccentricity bias and less unfairness in contexts with such implications. Overall, this corroborates the adequacy of EAUC and its potential applicability to guide the construction of fair models.
@@
RMSE MAE EAUC
Random 1.696 ± 0.000 1.386 ± 0.001 0.395 ± 0.004
Dyad Average 0.964 ± 0.000 0.781 ± 0.000 0.435 ± 0.008
MF 0.826 ± 0.001 0.635 ± 0.001 0.354 ± 0.006
w/ Linear corr. 1.186 ± 0.004 0.894 ± 0.003 0.253 ± 0.001
w/ Linear corr. + ML-RUS (Clipping) 0.879 ± 0.001 0.686 ± 0.001 0.307 ± 0.001
w/ Linear corr. + ML-RUS (Sigmoid) 0.977 ± 0.001 0.784 ± 0.002 0.267 ± 0.001
w/ Random Forest corr. 1.008 ± 0.001 0.809 ± 0.001 0.276 ± 0.001